OpenStudy (anonymous):

2x - y + z = -4 z = 5 -2x + 3y - z = -10

OpenStudy (agent0smith):

z=5, so replace all z's with 5 is your first step.

OpenStudy (anonymous):

2x - y + 5 = -4 z = 5 -2x + 3y - 5 = -10 ..

OpenStudy (agent0smith):

Now you can simplify the first and last equations a bit

OpenStudy (anonymous):

I'm really bad at simplifying.. x + 2y = -6?

OpenStudy (agent0smith):

I'm not sure how you got that... start with the first one: 2x - y + 5 = -4 subtract 5 on both sides

OpenStudy (anonymous):

2x - y = -9

OpenStudy (agent0smith):

Much better. Now try simplifying the third equation: -2x + 3y - 5 = -10

OpenStudy (anonymous):

-2x + 3y = -5

OpenStudy (agent0smith):

Good. So now we have z=5, and these two: 2x - y = -9 -2x + 3y = -5 hopefully you kinda know how to use elimination, cos it'll make it very easy

OpenStudy (anonymous):

Sorta. Substitution and elimination is what i'm learning right now.

OpenStudy (agent0smith):

Use elim. Add the two equations together, vertically, like so 2x - y = -9 -2x + 3y = -5 --------------